Set up the problem to be done to find the number asked for in the expression "six less than twice 29/6?
Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)
Answer by Alan3354(69443)

Set up the problem to be done to find the number asked for in the expression "six less than twice 29/6?
---------
n = 2*(29/6) - 6
= 29/3 - 6
= 29/3 - 18/3
= 11/3